OpenStudy (anonymous):

A fan club that comprises of boys and girls have a total of 160 members. if 32 more girls join the club, the percentage of girls will increase by 5%. Find the number of boys in the club.

myininaya (myininaya):

so from the first sentence we have b+g=160 do you see that like b is the number of boys and g is the number of girls

OpenStudy (anonymous):

and then: b+1.05g=160+32

OpenStudy (anonymous):

I don't get it??? Why 1.05g?

OpenStudy (anonymous):

because g gets increased by 5%, so 1+(5/100)

OpenStudy (anonymous):

oh okay...then how do I work it out from there???

OpenStudy (anonymous):

from the first equation you substract g from both sides, so: b= 160-g and then you substitute this value of b on the second equation: 160-g+1.05g=160+32 0.05g=32 g=32/0.05, are you sure its 5%, because the value of g, doesn't make sense....

OpenStudy (anonymous):

yeahh...but the answer is actually 48 boys. But i dunno hot to get it...